Write down pauli matrices.
1
Expert's answer
2020-11-02T09:26:42-0500

Solution pauli matrices can be written as


"\sigma_x=\begin{pmatrix}
 0 & 1 \\
 1 & 0
\end{pmatrix}"



"\sigma_y=\begin{pmatrix}
 0 & -i\\
 i & 0
\end{pmatrix}"

i is iota.

And


"\sigma_z=\begin{pmatrix}
 1 & 0\\
 0 & -1
\end{pmatrix}"